OpenStudy (anonymous):

find the equation of a line that is parallel to the line y=12x and contains the point(1,-3)

OpenStudy (luigi0210):

So you know the slope, right?

OpenStudy (anonymous):

it's 12 right?

OpenStudy (luigi0210):

Yup! Now you use the point-slope formula and the point (1, -3)

OpenStudy (anonymous):

slope is 12.and (x1,y1)=(1,-3) requird equation is, (y-y1)=m(x-x1) m is the slope.now put all the values u will get the equation